What could be reason why a pendulum clock was accurate in lower ground but needed to be adjusted when moved to higher ground
1
Expert's answer
2020-06-23T12:59:26-0400

In pendulum clock the timekeeping element is made in form of swinging weight. The rate of pendulum is strongly influenced by the local gravity (see https://en.wikipedia.org/wiki/Pendulum_clock#Local_gravity). Generally period is proportional to "\dfrac{1}{\sqrt{g}}" and the gravitational acceleration is proportional to "\dfrac{1}{r^2}," where "r" is the distance from the center of the Earth. Therefore, the period will be greater at the greater height and the clock should be adjusted. The same effect can be observed when moving the clock from polar latitudes to the equatorial regions.